The loop of the curve has an equation of y² = x (1 - x)^2. Find the



area enclosed by the loop of the curve.

1
Expert's answer
2022-06-08T13:45:34-0400

Simplify the function: "y=\pm \sqrt{x}(1-x)".

Limits of integration: "x=0", "x=1".




"S=2\int\limits_0^1\sqrt{x}(1-x)dx=2\int\limits_0^1(\sqrt{x}-x\sqrt{x})dx=2\int\limits_0^1(x^{1/2}-x^{3/2})dx="

"=2(\frac{x^{3/2}}{3/2}-\frac{x^{5/2}}{5/2} )|_0^1=2(\frac{2x^{3/2}}{3}-\frac{2x^{5/2}}{5} )|_0^1=2(\frac{2}{3}-\frac{2}{5})=\frac{8}{15}".